# NV-Embed: Improved Techniques for Training LLMs as Generalist Embedding Models

### v2

@@ Abstract
Decoder-only large language model (LLM)-based embedding models are beginning to outperform BERT or T5-based embedding models in general-purpose text embedding tasks, including dense vector-based retrieval. In this work, we introduce the NV-Embed model, incorporating architectural designs, training procedures, and curated datasets to significantly enhance the performance of LLM as a versatile embedding model, while maintaining its simplicity and reproducibility. For model architecture, we propose a latent attention layer to obtain pooled embeddings, which consistently improves retrieval and downstream task accuracy compared to mean pooling or using the last <EOS> token embedding from LLMs. To enhance representation learning, we remove the causal attention mask of LLMs during contrastive training. For training algorithm, we introduce a two-stage contrastive instruction-tuning method. It first applies contrastive training with instructions on retrieval datasets, utilizing in-batch negatives and curated hard negative examples. At stage-2, it blends various non-retrieval into instruction tuning, which not only enhances non-retrieval task accuracy but also improves retrieval performance. For training data, we utilize the hard-negative mining, synthetic data generation and existing public available datasets to boost the performance of embedding model. By combining these techniques, our NV-Embed-v1 and NV-Embed-v2 models obtained the No.1 position on the Massive Text Embedding Benchmark (MTEB) (as of May 24, 2024 and August 30, 2024, respectively) across 56 embedding tasks, demonstrating the sustained effectiveness of the proposed methods over time. Additionally, it achieved the highest scores in the Long Doc section and the second-highest scores in the QA section of the AIR Benchmark, which covers a range of out-of-domain information retrieval topics beyond those in MTEB. We open-source the model at: https://huggingface.co/nvidia/NV-Embed-v2.
@@ 1 Introduction
    For model architecture, we propose a novel latent attention layer to obtain pooled embeddings for a sequence of tokens. In contrast to the popular average pooling in bidirectional embedding models (e.g., Wang et al., 2022) and the last <EOS> token embedding in decoder-only LLMs (Neelakantan et al., 2022; Wang et al., 2023b), our proposed pooling technique consistently improves the accuracy of retrieval and other downstream tasks. To further enhance the representation learning, we remove the causal attention mask during the contrastive training of decoder-only LLM, resulting in solid improvements. Our design is simpler yet more effective compared to recent related work (BehnamGhader et al., 2024; Muennighoff et al., 2024), which involves an additional training phase with masked token prediction or a mixed training objective.
@@
We organize the rest of the paper in the following. In § 2, we discuss the related work. We present the architectural and training method in § 3. We provide detailed recipe of training data curation in § 4. We present the experiment results in § 5 and conclude the paper in § 6. AIR-bench results are shown in § A.
@@ 3.2 Latent Attention Layer
Figure 1: The illustration of proposed architecture design comprising of decoder-only LLM followed by latent attention layer. Latent attention layer functions as a form of cross-attention where the decoder-only LLM output serves as queries ($Q$) and trainable latent array passes through the key-value inputs, followed by MLP. Blue dotted lines indicate the two matrix multiplications involved in QKV-attentions.
@@ 4.1 Public Retrieval Datasets
It is important to note that certain datasets (e.g., MSMARCO) are training splits of the MTEB Benchmark, which we follow the existing practices established by leading generalist embedding models (Meng et al., 2024b; Wang et al., 2023b; BehnamGhader et al., 2024; Muennighoff et al., 2024). Table 9 further provides the number of samples used for training. We demonstrate the zero-shot generalization capability of NV-Embed on AIR-bench in A.
@@ 4.3 Synthetic Tasks Dataset
Due to the limited variety of subjects and tasks in public training datasets, the available instruction templates for training are also restricted. To enhance task-wise generalization, we employ the Mixtral-8x22B-Instruct-v0.1 model (MistralAI, ) to create a dataset consisting of 120,000 synthetic examples across 60,000 synthetic tasks. Following a two-step prompting approach proposed by E5-mistral-7b-instruct (Wang et al., 2023b), we adjust the prompts for Mixtral-8x22B-Instruct-v0.1 and English text. We generate only the short-long, long-short, and short-short examples (40,000 of each), as we use public STS datasets and do not assess bitext retrieval tasks. Example prompts for synthetic data generation can be found in Appendix 12 and 13.
@@ 5 Experiments
5.1 Experimental Details
We train Mistral 7B LLM model end-to-end with a contrastive loss using LoRA with rank 16, alpha 32 and dropout rate of 0.1. We use Adam optimizer with 50 warm-up steps and learning rate 2e-5 for first stage and 1.5e-5 for second stage with linear decay. The model is finetuned with 128 batch size, where each batch is composed of a query paired with 1 positive and 7 hard negative documents. Training samples from different datasets in Table 9 are uniformly sampled. We train using Bfloat16, and set the maximum sequence length as 512 tokens. The special <BOS> and <EOS> tokens are appended at the start and end of given query and documents. The whole training is conducted in two stages where the model is initially trained on retrieval datasets utilizing in-batch negative technique. Subsequently, the model is trained with blended datasets with both retrieval and non-retrieval embedding tasks.
@@ 5.2 MTEB Results
We evaluate the proposed NV-Embed model on the full MTEB benchmark (Muennighoff et al., 2022) across 56 tasks. Table 1 summarizes the averaged MTEB scores for seven sub-category tasks compared to the frontier models on the MTEB leaderboard [^2] https://github.com/embeddings-benchmark/mteb. Our initial model, namely NV-Embed-v1 get the score of 69.32 and obtain the No.1 position on the MTEB as of May 24, 2024 (detailed benchmark scores available in Table 2). We then further improve the model through the curation of training dataset, including adding more retrieval datasets, applying positive-aware hard-negative mining technique, using synthetic data generation process and constructing example-based multi-class labels. As a result, our NV-Embed-v2 model sets a new record high score of 72.31 and reclaimed No.1 (as of Aug 30, 2024) on the highly competitive MTEB leaderboard, further highlighting the sustained effectiveness of the proposed methods. In following sub-section 5.3, we will present the ablation studies on the design choices regarding the model architecture, training algorithm and the curation of training data.
@@
SFR-Embedding-2R (Meng et al., 2024b) demonstrates competitive scores on the MTEB (70.31) and BEIR (60.18) benchmarks by continuing to finetune the e5-mistral-7b-instruct model (Wang et al., 2023b). However, it remains largely constrained by the architectural limitations of its parent model, such as the causal attention mask and the last token pooling method. In contrast, our NV-Embed model is trained starting from the Mistral 7B LLM (Jiang et al., 2023) rather than finetuning e5-mistral-7b-instruct (Wang et al., 2023b). It features a new architecture that removes the unnecessary causal attention mask and further improves the sequence pooling mechanism with a latent attention layer. Table 3 and 11 provides a detailed scores of BEIR and MTEB benchmarks.
@@ 6 Conclusion
We introduced the NV-Embed model, a decoder-only LLM designed to outperform existing bidirectional models in general-purpose text embedding tasks. For model architecture, we propose a latent attention layer to obtain expressive pooled embeddings and remove the unnecessary causal attention mask of decoder-only LLMs. For training algorithm, we introduce a two-stage contrastive instruction-tuning scheme to sequentially improve the embedding tasks. By leveraging carefully curated datasets, hard-negative mining, synthetic data generation and example-based multi-class labeling, our approach achieve the superior accuracy across diverse embedding tasks. As a result, the series of NV-Embed models achieved and maintained the No.1 ranking on the MTEB leaderboard and also demonstrated superior accuracy in out-of-domain tasks in AIR Benchmark. The sustained effectiveness of NV-Embed highlights the importance of our proposed architecutre design, training procedures and dataset curation in achieving state-of-the-art performance in the evolving landscape of text embedding models.
@@ 7 Acknowledgment
Appendix A Air Benchmark
@@
In this first appendix section, we present AIR-Bench [^3] https://github.com/AIR-Bench/AIR-Bench (version of 24.04) that is newly released information retrieval benchmark, incorporating the diverse and comprehensive domains such as healthcare, law, news, book, arxiv, finance and synthetically generated samples using diverse LLMs. Importantly, AIR-Bench can help us to understand the generalization capability of the embedding/retrieval model, because the majority of different domain samples do not appear in MTEB benchmarks. Moreover, the AIR-Bench is designed as a closed-book benchmark whose ground truth is kept confidential. As a result, the benchmark score can be only obtained through the HuggingFace Hub platform.
@@
Appendix B Experimental Details and Instruction Templates for Training and Evaluation
We use the Adam optimizer for each training stage. The optimizer hyperparameters are included in Table 8. We restart the optimizer with the same 50 warm-up steps and lower learning rate for the second stage.
@@ Appendix B Experimental Details and Instruction Templates for Training and Evaluation
Appendix C Latent-Attention Visualization
@@
Touche2020 30.26 30.57 28.18 29.94 29.67 28.38 88.44
TREC-COVID 64.27 82.26 87.28 85.98 78.11 85.88 31.78

### v3

@@ Abstract
Decoder-only large language model (LLM)-based embedding models are beginning to outperform BERT or T5-based embedding models in general-purpose text embedding tasks, including dense vector-based retrieval. In this work, we introduce the NV-Embed model, incorporating architectural designs, training procedures, and curated datasets to significantly enhance the performance of LLM as a versatile embedding model, while maintaining its simplicity and reproducibility. For model architecture, we propose a latent attention layer to obtain pooled embeddings, which consistently improves retrieval and downstream task accuracy compared to mean pooling or using the last <EOS> token embedding from LLMs. To enhance representation learning, we remove the causal attention mask of LLMs during contrastive training. For training algorithm, we introduce a two-stage contrastive instruction-tuning method. It first applies contrastive training with instructions on retrieval datasets, utilizing in-batch negatives and curated hard negative examples. At stage-2, it blends various non-retrieval into instruction tuning, which not only enhances non-retrieval task accuracy but also improves retrieval performance. For training data, we utilize the hard-negative mining, synthetic data generation and existing public available datasets to boost the performance of embedding model. By combining these techniques, our NV-Embed-v1 and NV-Embed-v2 models obtained the No.1 position on the Massive Text Embedding Benchmark (MTEB) (as of May 24, 2024 and August 30, 2024, respectively) across 56 embedding tasks, demonstrating the sustained effectiveness of the proposed methods over time. Also, it achieved the highest scores in the Long Doc section and the second-highest scores in the QA section of the AIR Benchmark, which covers a range of out-of-domain information retrieval topics beyond those in MTEB. We further provide the analysis of model compression techniques for generalist embedding models. We open-source the model at: https://huggingface.co/nvidia/NV-Embed-v2.
@@ 1 Introduction
    For model architecture, we propose a novel latent attention layer to obtain pooled embeddings for a sequence of tokens. In contrast to the popular average pooling in bidirectional embedding models (e.g., Wang et al., 2022) and last <EOS> token embedding in decoder-only LLMs (Neelakantan et al., 2022; Wang et al., 2023b), our proposed pooling technique consistently improves accuracy of retrieval and other downstream tasks. To further enhance representation learning, we remove causal attention mask during contrastive training of decoder-only LLM, resulting in solid improvements. Our design is simpler yet more effective compared to related work (BehnamGhader et al., 2024; Muennighoff et al., 2024), which involves an additional training phase with masked token prediction or a mixed training objective.
@@
  5. 5.
    We study the model compression techniques, including pruning, quantization and knowledge-distillation, for LLM-based embedding models. Through the comparison with smaller embedding models directly built on Llama3.2-3B, Qwen2.5-3B, and Minitron-4B, we demonstrate that our model compression approach achieves superior accuracy and quantization robustness.
We organize the rest of the paper in the following. In § 2, we discuss the related work. We present the architectural and training method in § 3. We provide detailed recipe of training data curation in § 4. We present the experiment results in § 5 and conclude the paper in § 6. Model compression techniques and results are presented in § A due to the page limit. AIR-bench results are shown in § B.
@@ 3.2 Latent Attention Layer
Figure 1: Proposed architecture design comprising of decoder-only LLM followed by latent attention layer. Latent attention layer functions as a form of cross-attention where the decoder-only LLM output serves as queries ($Q$) and trainable latent array passes through the key-value inputs, followed by MLP. Blue dotted lines indicate the two matrix multiplications involved in QKV-attentions.
@@ 4.1 Public Retrieval Datasets
It is important to note that certain datasets (e.g., MSMARCO) are training splits of the MTEB Benchmark, which we follow the existing practices established by leading generalist embedding models (Meng et al., 2024b; Wang et al., 2023b; BehnamGhader et al., 2024; Muennighoff et al., 2024). Table 12 further provides the number of samples used for training. We demonstrate the zero-shot generalization capability of NV-Embed on AIR-bench in B.
@@ 4.3 Synthetic Tasks Dataset
Due to the limited variety of subjects and tasks in public training datasets, the available instruction templates for training are also restricted. To enhance task-wise generalization, we employ the Mixtral-8x22B-Instruct-v0.1 model (MistralAI, ) to create a dataset consisting of 120,000 synthetic examples across 60,000 synthetic tasks. Following a two-step prompting approach proposed by E5-mistral-7b-instruct (Wang et al., 2023b), we adjust the prompts for Mixtral-8x22B-Instruct-v0.1 and English text. We generate only the short-long, long-short, and short-short examples (40,000 of each), as we use public STS datasets and do not assess bitext retrieval tasks. Example prompts for synthetic data generation can be found in Appendix 15 and 16.
@@ 5 Experiments
Training and inference experiment details are illustrated in Appendix C.
@@ 5.1 MTEB Results
We evaluate the proposed NV-Embed model on the full MTEB benchmark (Muennighoff et al., 2022) across 56 tasks. Table 1 summarizes averaged MTEB scores for seven sub-category tasks compared to frontier models on MTEB leaderboard [^2] https://github.com/embeddings-benchmark/mteb. Our initial model, namely NV-Embed-v1 get the score of 69.32 and obtain the No.1 position on the MTEB as of May 24, 2024 (detailed benchmark scores available in Table 2). We then further improve the model through the curation of training dataset, including adding more retrieval datasets, applying positive-aware hard-negative mining technique, using synthetic data generation process and constructing example-based multi-class labels. As a result, our NV-Embed-v2 model sets a new record high score of 72.31 and reclaimed No.1 (as of Aug 30, 2024) on highly competitive MTEB leaderboard, further highlighting the sustained effectiveness of the proposed methods. In following sub-section 5.2, we will present ablation studies on design choices regarding the model architecture, training algorithm and the curation of training data.
@@
SFR-Embedding-2R (Meng et al., 2024b) demonstrates competitive scores on the MTEB (70.31) and BEIR (60.18) benchmarks by continuing to finetune the e5-mistral-7b-instruct model (Wang et al., 2023b). However, it remains largely constrained by the architectural limitations of its parent model, such as the causal attention mask and the last token pooling method. In contrast, our NV-Embed model is trained starting from the Mistral 7B LLM (Jiang et al., 2023) rather than finetuning e5-mistral-7b-instruct (Wang et al., 2023b). It features a new architecture that removes the unnecessary causal attention mask and further improves the sequence pooling mechanism with a latent attention layer. Table 3 and 14 provides a detailed scores of BEIR and MTEB benchmarks.
@@ 6 Conclusion
We introduced the NV-Embed model, a decoder-only LLM designed to outperform existing bidirectional models in general-purpose text embedding tasks. For model architecture, we propose a latent attention layer to obtain expressive pooled embeddings and remove the unnecessary causal attention mask of decoder-only LLMs. For training algorithm, we introduce a two-stage contrastive instruction-tuning scheme to sequentially improve the embedding tasks. By leveraging carefully curated datasets, hard-negative mining, synthetic data generation and example-based multi-class labeling, our approach achieve the superior accuracy across diverse embedding tasks. As a result, the series of NV-Embed models achieved and maintained the No.1 ranking on the MTEB leaderboard and also demonstrated superior accuracy in out-of-domain tasks in AIR Benchmark.
@@ 7 Acknowledgment
Appendix A Comprehensive Study of Model Compression Techniques for NV-Embed
Increasing computational and memory demands of LLM-based embedding model present the challenges for the deployment, limiting their scalability and accessibility. In this appendix section, we provide the analysis of post-training model compression techniques (i.e., pruning and quantization) for generalist embedding models. Our analysis demonstrates that these compression methods enhance the accuracy and robustness of LLM-based embedding models, surpassing the performance of smaller-sized embedding models based on Llama3.2-3B, Qwen2.5-3B and Minitron-4B.
In model compression process, we first perform pruning the NV-Embed-v2 model, reducing its size from 8 billion parameters to 3.5 billion (i.e., pruning the main decoder-only blocks and removing the latent attention block). Next, we apply quantization to lower its precision to 8-bit weights including integer and floating (E4M3, E5M2) formats. Finally, we perform continual re-training using fine-tuning (PEFT) method known as low-rank adaptation (LoRA) to restore the model’s accuracy. For evaluation, we evaluate our model on MTEB benchmark (Muennighoff et al., 2022).
A.1 Pruning
In order to find better pruning techniques, we apply three methods (magnitude-based, WANDA(Sun et al., 2023), SparseGPT(Frantar & Alistarh, 2023)) for semi-structured (2:4 and 4:8) and unstructured approaches. Note, unstructured pruning strategy removes the network elements from individual weights, while the structured strategy removes the blocks of nonzero weights in higher granularity ways such as row/columns of weight metrics. Semi-structured is the hardware friendly way (N:M sparsity), ensuring that N weights remain non-zero within every group of M weights. For example, 4:8 semi-structured pruning prunes four out of every eight elements in a weight tensor. This semi-structured sparsity reduces the size of the weight matrices and computational cost, while maintaining certain regularity for efficient hardware utilization. The literature presents various criteria for determining which weights to prune. The simplest approach is magnitude-based pruning, which retains weights with higher absolute values and removes the rest. Another approach is WANDA (Sun et al., 2023) which introduces a pruning technique that considers both weights and activations. SparseGPT (Frantar & Alistarh, 2023) identifies the non-critical connections by utilizing the approximate hessian based optimization method.
Table 6 summarizes the averaged MTEB scores for different model pruning, respectively. Among these techniques, SparseGPT generally delivers the best results, while magnitude-based and WANDA methods produce comparable performance both during pruning and after retraining as shown in Table 6. Notably, semi-structured (2:4) pruning yields the lowest scores but demonstrates the greatest accuracy recovery following retraining for MTEB benchmarks. Based on these findings, we focus on SparseGPT pruning for subsequent ablation studies.
Table 6: Pruning - MTEB benchmark
Pruning Criterion Semi-structured Unstructured
2:4 4:8
Magnitude Pruning 64.62 67.6 69.18
Re-train 69.96 70.46 70.84
Wanda Pruning 64.26 67.87 70.19
Re-train 69.74 70.42 70.81
SparseGPT Pruning 68.48 70.11 71.33
Re-train 70.41 70.9 71.18
A.2 Knowledge distillation
In traditional accuracy recovery approaches after model compression, ground truth labels are utilized for continual retraining. To improve this retraining process, we leverage a knowledge distillation loss term, where the uncompressed model serves as the teacher, transfering the knowledge of the more advanced teacher model to a smaller and simpler student model. To enable the student model mimic the teacher’s behavior, we introduce mean-squared error losses for both the output state ($S_{o}$) and the intermediate states ($S_{1}-S_{o-1}$).
For this knowledge distillation process, we use the the uncompressed embedding model serves as the teacher, while the compressed version acts as the student. We remove the latent attention block and compensate the accuracy degradation with knowledge distillation. The knowledge distillation loss is defined as $L_{kd}=\sum_{n=1}^{O-2}[MSE(S_{s}^{n},S_{t}^{n})]+MSE(S_{s}^{O-1},S_{t}^{O})$ where $L_{kd}$ is knowledge distillation loss, O is the number of layers, n is layer number, MSE represents the mean-squared function, $S_{s}$ is student state and $S_{t}$ is the teacher state. Based on this, the total loss function is sum of contrastive and knowledge distillation loss as: $L_{total}=L_{contrastive}+\alpha\times L_{kd}$ where $\alpha$ is weight term.
As presented in Table 7, incorporating knowledge distillation ("GT+KD") consistently outperforms using only ground truth labels ("GT") across different approaches for MTEB benchmarks. Among the methods, 2:4 semi-structured pruning yields the worst results but benefits the most from knowledge distillation, achieving improvements of 0.76 on the MTEB benchmark.
Table 7: Knowledge Distillation - MTEB benchmark
Label Types Semi-structured Unstructured
2:4 4:8
GT 70.41 70.90 71.18
GT+KD 71.17 71.22 71.48
A.3 Quantization
For weight quantization stage, we adopt GPTQ (Frantar et al., 2022), a post-training weight quantization method that utilizes approximate Hessian information to reduce the precision of the weights. To evaluate our compressed embedding models, we compare them against three smaller LLM-based embedding models—Llama3.2-3B, Qwen2.5-3B, and Minitron-4B—which have varying numbers of weight parameters. Table 8 provides the averaged MTEB scores for compressed models (pruning and quantization), respectively.
A key observation is that our compressed models demonstrates superior robustness in low precision settings compared to their smaller counter parts.For example, NV-Embed quantized to INT8 maintains nearly identical MTEB scores (0.0% for 2:4 semi-structured, 0.01% for 4:8 semi-structured, and 0.01% for unstructured) compared to the performance drops observed in smaller models such as Llama-3B (-0.47%), Qwen-3B (-0.14%), and Minitron-4B (-0.84%). This trend remains consistent across different 8 bit precision cases as well.
Compared to integer format which has an uniform numerical distribution, floating point format can also represent the same number of discrete points, covering larger numerical range and non-uniform distributions (high precision for small values and lower precision for large values). There are two primary FP8 format: E4M3 (4-bit exponent, 3-bit mantissa), E5M2 (5-bit exponent, 2-bit mantissa) where 1 bit represents the signed bit. Table 8 shows that 8 bit floating point (E4M3 and E5M2) achieve comparable MTEB scores to the INT8 format.
Table 8: Quantization - MTEB benchmark
Precision 16bit INT8 FP8 (E4M3) FP8 (E5M2)
NV-Embed (2:4) Score 71.17 71.17 70.94 71.14
Diff (%) - $0$34% 0.03%
NV-Embed (4:8) Score 71.22 71.23 71.28 71.48
Diff (%) - 0.01% 0.08% 0.37%
NV-Embed (Unstr) Score 71.48 71.49 71.55 71.75
Diff (%) - 0.01% 0.09% 0.37%
Llama3.2-3b Score 70.31 69.98 70.05 70.06
Diff (%) - -0.47% -0.36% -0.35%
Qwen2.5-3b Score 69.77 69.70 69.70 69.67
Diff (%) - -0.1% -0.1% -0.14%
Minitron-4b Score 70.68 70.09 69.97 69.97
Diff (%) - -0.84% -1.0% -1.02%
Appendix B Air Benchmark
@@
In this appendix section, we present AIR-Bench [^3] https://github.com/AIR-Bench/AIR-Bench (version of 24.04) that is newly released information retrieval benchmark, incorporating the diverse and comprehensive domains such as healthcare, law, news, book, arxiv, finance and synthetically generated samples using diverse LLMs. Importantly, AIR-Bench can help us to understand the generalization capability of the embedding/retrieval model, because the majority of different domain samples do not appear in MTEB benchmarks. Moreover, the AIR-Bench is designed as a closed-book benchmark whose ground truth is kept confidential. As a result, the benchmark score can be only obtained through the HuggingFace Hub platform.
@@
Appendix C Experimental Details and Instruction Templates for Training and Evaluation
We train Mistral 7B LLM model end-to-end with a contrastive loss using LoRA with rank 16, alpha 32 and dropout rate of 0.1. We use Adam optimizer with 50 warm-up steps and learning rate 2e-5 for first stage and 1.5e-5 for second stage with linear decay. The optimizer hyperparameters are included in Table 11. We restart the optimizer with the same 50 warm-up steps and lower learning rate for the second stage. The model is finetuned with 128 batch size, where each batch is composed of a query paired with 1 positive and 7 hard negative documents. Training samples from different datasets in Table 12 are uniformly sampled. We train using Bfloat16, and set the maximum sequence length as 512 tokens. The special <BOS> and <EOS> tokens are appended at the start and end of given query and documents. The whole training is conducted in two stages where the model is initially trained on retrieval datasets utilizing in-batch negative technique. Subsequently, the model is trained with blended datasets with both retrieval and non-retrieval embedding tasks.
@@ Appendix C Experimental Details and Instruction Templates for Training and Evaluation
Appendix D Latent-Attention Visualization
@@
Touche2020 30.26 30.57 28.18 29.94 29.67 28.38 31.78
TREC-COVID 64.27 82.26 87.28 85.98 78.11 85.88 88.44